OpenStudy (anonymous):

For numbers 1 – 5: A. Write whether the problem is asking you to find SURFACE AREA or VOLUME B. Write the formula which you would use to solve the problem C. Solve and show all your work D. Label your answer with the correct units 1. Sam is in charge of filling a rectangular hole with cement. The hole is 9 feet long, 3 feet wide and 2 feet deep. How much cement will Sam need? 2. The cylindrical canister of a fire extinguisher has a radius of 4 inches and is 12 inches high. How many exact cubic inches can it hold?

OpenStudy (anonymous):

Oh gosh geometry okay give me a second

OpenStudy (anonymous):

Yeah I Hate it

OpenStudy (anonymous):

For 1. a. It would definitely be volume because if you fill a hole with cement you need to fill the whole thing not just the outside of it.

OpenStudy (anonymous):

1. b would be Volume=Length*width*Height

OpenStudy (anonymous):

O OKay

OpenStudy (anonymous):

1.c would be Volume=9*3*2=54

OpenStudy (anonymous):

then 1. d. would be 54 Ft.^3 (cubed)

OpenStudy (anonymous):

can i send you the other questions as well if its not to much trouble

OpenStudy (anonymous):

sure do you want help with two as well?

OpenStudy (anonymous):

Yes Please

OpenStudy (anonymous):

This is whats do

OpenStudy (anonymous):

due

OpenStudy (anonymous):

Okay so for number 2. a is volume for 2 reasons because it talks about cubic inches which means volume and its asking how much can it hold

OpenStudy (anonymous):

okay got it

OpenStudy (anonymous):

2b. Volume=(Pie)R^2*h

OpenStudy (anonymous):

2.c. volume= (3.14)(4)^2(12)=602.88 and 2.D. would be 602.88 inches

OpenStudy (anonymous):

Okay it starting to make sense Now

OpenStudy (anonymous):

This is #3 3. Jessica is decorating conical party hats for her party by wrapping them in colored tissue paper. Each hat has a radius of 4 centimeters and a slant height of 8 centimeters. If she wants to wrap 6 party hats, how much exact paper will she need?

OpenStudy (anonymous):

megsk133u still there

OpenStudy (anonymous):

@creegan

OpenStudy (anonymous):

This is #3 3. Jessica is decorating conical party hats for her party by wrapping them in colored tissue paper. Each hat has a radius of 4 centimeters and a slant height of 8 centimeters. If she wants to wrap 6 party hats, how much exact paper will she need?

OpenStudy (anonymous):

192cm

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

yes because if u have length times height you do 8*4=32*6=192cm

OpenStudy (anonymous):

Okay I get it

OpenStudy (anonymous):

yup